They needed $486,000 to start the business. They invested in the ratio of 13:14. How much money did Clinton invest?
Answer by Alan3354(31487)

They needed $486,000 to start the business. They invested in the ratio of 13:14. How much money did Clinton invest?
-------------
13+14 = 27
486000/27 = 18000
-------
13*18000 = 234000
14*18000 = 252000